OpenStudy (anonymous):

which ordered pair is the solution to this equation? -2x+9y=-26 a.(-4,-4) b.(4,4) c.(-4,-5) d.(-5,-4)

OpenStudy (smurfy14):

here is the graph of the line

OpenStudy (anonymous):

okhay whats the answer?

OpenStudy (smurfy14):

well none of thoses are on the line so are you sure you typed the answers right? if you did then i think there is a typo

OpenStudy (anonymous):

thats it... thats how it was

OpenStudy (anonymous):

written

OpenStudy (radar):

Try d. -2(-5) + 9(-4) =10-36=-26 Looks like it would work.

OpenStudy (anonymous):

radar got the right answer, d. A solution using Mathematica with comments is attached.

OpenStudy (radar):

Interesting pdf. I like the fact it checked the wrong answers also producing false outputs.

OpenStudy (anonymous):

radar, Thank you.